OpenStudy (anonymous):

May I please have help with algebra one problems, I have to write an equation in point-slope form for the line through the given point, (4,9), and (-2,-6)

OpenStudy (anonymous):

y-y1=m(x-x1) m=(y2-y1)/(x2-x1)

OpenStudy (anonymous):

(x1,y1)=(4,9)

OpenStudy (anonymous):

thank you! I greatly appreciate it!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

What's that supposed to mean?

OpenStudy (anonymous):

you are welcome.

OpenStudy (anonymous):

Oh haha